The denominator of a fraction exceeds the numerator by 2. If 5 be added to numerator the fraction increased by unity. The Fraction is?
Answer by mananth(16946)

numerator be x
denominator = x+2
fraction = x/(x+2)
..
(x+5)/(x+2) = x/(x+2) +1
..
(x+5)/(x+2) - x/(x+2) =1
(x+5 -x)/(x+2)= 1
5/(x+2) = 1
5 = x+2
5-2 = x
x=3
...
The fraction is 3/5
.
m.ananth@hotmail.ca